HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE MARVELOUS WORLD OF OH^{+}

\underline{MARCO PEZZELLA} , SERGEI N. YURCHENKO, JONATHAN TENNYSON, Department of Physics and Astronomy, University College London,

London WC1E 6BT, UK

The OH^{+} ion is an interesting molecular species, due to its critical intermediate in interstellar chemistry. OH^{+}, together with H_{2}O^{+} and H_{3}O^{+}, is an important probe of the cosmic ray ionization rates in diffuse clouds, on the surfaces of molecular clouds. It is formed by the reaction of oxygen atoms with cosmic ray ionised H or H_{2}, and it goes to further reactions with other hydrogen sources to form, via multiple steps, OH^{1}. It acts as an enabler of more complex reaction networks occurring in the interstellar medium^{2}. The molecule is also formed in the ionosphere of Earth and comets^{3}. We started by collecting transition from different experimental work^{4} and processing them using MARVEL^{5} for generating high level accuracy energy levels, with an uncertainty of 1 × 10^{−2} cm^{−1}. These energies are used in DUO^{6} to fit the potential energy curve and couplings of the X^{3}Σ state, leading to a new linelist with RMSE=5.43×10^{−2} cm^{−1} with respect to experimental transition energies.

The same procedure is applied to the A^{3}Π state. The coupling between this state and the a^{1}Δ, b^{1}Σ^{+}, and c^{1}Π, leads to a more complex situation to analyse^{7}. The

1D. Hollenbach et al 2012 ApJ 754 105

2A. J. Porras et al 2014 ApJL 781 L8

3R. Martinez et al 2005 JCP 123 174312, P.V. Stoeva et al 2005 Planetary and Space Science 53 327

4Rehfuss et al 1992 JMS 151 59,C. R. Markus et al 2016 ApJ 817 138,J. N. Hodges et al 2017 ApJ 840 81

5T. Furtenbacher et al 2007 JMS 245 115

6Yurchenko et al 2016 CPC 202 262

7D. Yarkony 1993 JPC 97 111 p-number: p151  Submitted on Wed, 14 Jun 2023 19:26:43 +0200

HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

X^{3}Σ and A^{3}Π state curves will be used to generate a comprehensive line list for OH^{+}.

p-number: p151  Submitted on Wed, 14 Jun 2023 19:26:43 +0200